Please help I do not understand.
Mahesha has $100 bills altogether in ten dollar bills and $5 bills
A. complete the table of values to show at least 4 possible combinations of $10 bills and $5 bills. Mahesha may have.
The table of values is like this
Number of 10$ bills
Number of 5$ bills
B. Is it possible for Mahesha to have eleven $5 bills. Explain.

Answer by josgarithmetic(39617)

x five dollar bills
y ten dollar bills
5x+10y=100
reduce or simplify
x+2y=50
Solve for y
2y=-x+50
y=-x/2+25

You should only pick natural number values for x and evaluate y. You can include 0 for x.

x=11 will not work. Try it and see.